SHARE VALUATION

QUESTION 8

Tobias Limited has issued an unlimited number of non-redeemable preference shares

with a par value of R100 each at a 4 percent (4%) discount. The preference dividend is

15 percent (15%) per year.

Management expects that the earnings per ordinary share will show no growth over the

next two years where-after it will grow at 6 percent (6%) infinitely. The most recent

dividend per share is R2,00.

The required rate of return of Tobias Limited is 20 percent (20%).

Required:

Calculate the value of the ordinary shares and preference shares of Tobias Limited.

(Show all necessary calculations).
Answer #1

1. Value of Preference Shares = Dividend / Rate = 15/(20%) = 75

2. Value of Ordinary Shares = 1357.14

Particular Year 0 (Present Value) Year 1 Year 2 Year 3 Discounting Factor @ 20%
Dividend 166.67       200.00               0.8333
Dividend 138.89       200.00               0.6944
Value 1051.59    1,514.29       212.00               0.6944
Total 1357.14

Value of Ordinary shares at the end of year 2 with growth rate (g) of 6% and cost of equity (ke) as 20%Dividend for year 1 and year 2 is discounted to its present value.

= D1/(Ke-g) = (200*1.06)/(20%-6%) = 1514.29

Value of Ordinary at the end of year 2 is discounted to its present value.